OpenStudy (anonymous):

indefinite integral (e^(sqrt(6x+4)))/(sqrt(6x+4))dx

OpenStudy (anonymous):

assume z=sqrt(6x+4) then dz =6/(2* sqrt(6x+4)) dx =3/sqrt(6x+4) dx thus the question redusec to integral of 1/3 * e^z dz and its integral is 1/3 * e^z +c = 1/3 * e^sqrt(6x+4) +c

OpenStudy (anonymous):

how did you get dz?

OpenStudy (anonymous):

I get it now. Thanks!

OpenStudy (anonymous):

welcome dear